OpenStudy (anonymous):

can anybody help with algebra 1

OpenStudy (anonymous):

what is your question?

OpenStudy (anonymous):

i sent the link its on page 280 task 3

OpenStudy (anonymous):

@master43 do you have algebra 1in CA

OpenStudy (anonymous):

the length of the necklace is the "independent variable" it determines the length of the string

OpenStudy (anonymous):

10 inches of necklace requires 200 inches of string, and each additional inch of necklace requires two additional inches of string making it linear

OpenStudy (anonymous):

so a 15 inch necklace should require 210 inches of string

OpenStudy (anonymous):

thank you and witch one is a b c or d cuze it has four steps to do?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i was looking at page 280 task 3 right?

OpenStudy (anonymous):

yes you where and they are asking to do 4 things that are marked as a b c d

OpenStudy (anonymous):

ok lets do them one by one

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

a) identify the "independent" and "dependent" variable the "independent variable" is the length of the necklace the "dependent variable" is the amount of string used

OpenStudy (anonymous):

ow ok and the rest

OpenStudy (anonymous):

b) write the graph as a "function rule" you could say in english "ten inches of necklace requires 200 inches of string, and each additional inch of necklace takes two more inches of string" or, you could say if \(n\) is then inches of necklace then the mount of string used is \[2n+180\]

OpenStudy (anonymous):

c) if you just plot the points given, the graph is "discrete" but if you connect the dots, it is "continuous"

OpenStudy (anonymous):

d) for a 15 inch necklace you would need \(2\times 15+180=30+180=210\) inches of string

OpenStudy (anonymous):

omg thank you so much

OpenStudy (anonymous):

yw